# DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference

### v1

@@ 3.4 Training and Inference
The loss component $\mathcal{L}_{K}$ corresponds to the cross-entropy loss calculated from the predictions of $\Psi(\zeta_{\boldsymbol{a}},\zeta_{\boldsymbol{r^{{}^{\prime}}}},\zeta_{\boldsymbol{k}})$, as defined in Eq. (14). Similarly, the aspect-only and debiased review-only losses are denoted as $\mathcal{L}_{A}$ and $\mathcal{L}_{R}$ respectively.
@@ 4 Experiments
Table 1: We retrained ${}^{\dagger}$, ${}^{\dagger}$ as fair baselines ensuring that comparisons are made under similar training settings, which is crucial for validating DINER’s superior performance.
@@ 4.1 Datasets
We conduct experiments on the ARTS dataset, as introduced by  Xing et al. (2020), to assess the efficacy of the proposed method. Detailed information about the ARTS dataset is shown in Appendix 5.
@@ 4.2 Baselines
Methods Laptop Restaurant
Vanill 74.96 79.26
$R\rightarrow L$ branch
 + Causal Intervention(NWGM) 75.44 81.02
 + Causal Intervention(IPW) 75.50 81.19
  +Counterfactual Reasoning 75.92 81.78
$A\rightarrow L$ branch
 + Counterfactual Inference 75.23 80.51
DINNER 76.51 82.46
RevTgt RevNon AddDiff Original
Laptop Vanill 62.45 85.93 76.33 80.41
DINNER 65.02($\uparrow$ 4.12%) 86.67($\uparrow$ 0.86%) 78.06($\uparrow$ 2.27%) 81.19($\uparrow$ 0.97%)
Restaurant Vanill 64.06 82.66 83.48 85.18
DINNER 70.69($\uparrow$ 10.35%) 83.56($\uparrow$ 1.08%) 86.07($\uparrow$ 3.10%) 87.32($\uparrow$ 2.51%)
Table 3: We use RoBERTa as the backbone. Vanill refers to ${}^{\dagger}$ in Table 1. We compare the Acc. on the Vanill and our DINNER framework. We also calculate the change of accuracy.
@@ 5 Result and Analysis
Table 1 presents a detailed comparison of various models’ performance for ARTS’laptop and restaurant domains, focusing on three key metrics: Acc., F1-score, and ARS.
@@
We evaluate DINER in two backbones: BERT-based and RoBERTa-based. These configurations are set to evaluate the effectiveness of DINER when integrated with different encoder-only PLMs. And DINER based on RoBERTa tends to outperform its BERT counterparts, which may be attributed to RoBERTa’s more robust pre-training on a larger and more diverse corpus, leading to better generalization capabilities Liu et al. (2019b). The results are compelling, showing that DINER(RoBERTa-based) model achieves the state-of-the-art performance across all metrics in both the laptop and restaurant domains, with a notable Acc. of 76.51 and 82.46, F1-scores of 73.27 and 76.92, and ARS of 59.40 and 64.02, respectively. DINER(RoBERTa-based) demonstrates superior performance in the Laptop domain, outpacing the baseline ${}^{\dagger}$ by margins of 1.55, 1.11, and 3.13 in terms of Acc., F1-score, and ARS metrics, respectively. In the Restaurant domain, the model further extends its lead, achieving improvements of 3.20, 6.45, and 4.06 in the same metrics. Similarly, the DINER(BERT-based) exhibits empirical enhancements.
@@ 5.1 More Detailed Result
We list in detail the performance of each model on the aforementioned three subsets of the ARTS dataset in Table 3. In the Laptop domain, the baseline model ${}^{\dagger}$ exhibits a RevTgt accuracy of 62.45, as RevTgt is the most challenging subset. It requires the model to pay precise attention to the target sentiment words. In contrast, the DINNER framework significantly enhances this metric to 65.02, marking a 4.12% increment. Similarly, for RevNon and AddDiff, DINNER outperforms the Vanilla baseline with modest improvements of 0.86% and 2.27%, respectively. The Restaurant domain further underscores the efficacy of the DINNER framework, where a remarkable 10.35% improvement is observed in the RevTgt task, elevating the accuracy from 64.06 to 70.69. The framework also exhibits gains in RevNon and AddDiff tasks by 1.08% and 3.10%, respectively. The significant improvement observed in the restaurant domain underscores the effectiveness of our methods, particularly given the inherently challenging nature of the test set data in this domain, as highlighted by Xing et al. (2020). Specifically, the more challenging the dataset, the greater the improvement our framework offers. Interestingly, our method also gives a slight improvement on the Original test set, illustrating the fact that we have also de-biased the robust test data on the Original test set.
@@
5.2 Effects of the two branches in DINER
For the $R\rightarrow L$ branch, NWGM Xu et al. (2015) yields a marginal improvement in accuracy across both domains. The method of IPW Pearl (2009) further enhances performance, suggesting the efficacy of backdoor adjustment intervention in the methods, and IPW has a more precise approximation compared to NWGM Xu et al. (2015). Due to DINER framework using hybrid intervention, IPW method can do Counterfactual Reasoning as described in Section 3.2, and performance is further enhanced upon the application of Counterfactual Reasoning.
5.3 Impact of Different Fusion Strategies
@@
$\left\{\begin{array}[]{l}\text{{MUL}-{Vanill}}:L_{a,r,x}=\zeta_{a}\cdot\zeta_{r}\cdot\zeta_{x},\\ \text{{MUL}-{sigmoid}}:L_{a,r,x}=\zeta_{x}\cdot\sigma(\zeta_{a})\cdot\sigma(\zeta_{r}),\\ \text{{MUL}-$\tanh$}:L_{a,r,x}=\zeta_{x}\cdot\tanh(\zeta_{a})\cdot\tanh(\zeta_{r}),\\ \text{{SUM}-{Vanill}}:L_{a,r,x}=\zeta_{a}+\zeta_{r}+\zeta_{x},\\ \text{{SUM}-{sigmoid}}:L_{a,r,x}=\zeta_{x}+\sigma(\zeta_{a})+\sigma(\zeta_{r}),\\ \text{{SUM}-$\tanh$}:L_{a,r,x}=\zeta_{x}+\tanh(\zeta_{a})+\tanh(\zeta_{r})\end{array}\right.$ (16)
The Acc. performance of six distinct different fusion strategies are reported in Table 4. From the table, we can find that the MUL fusion, regardless of the activation function, consistently underperforms in comparison to its SUM counterparts. Apparently, SUM fusion strategies are more stable and robust and more suitable for ASBA task. The superior performance of SUM fusion strategies, particularly with the $\tanh$ activation, underscores the effectiveness of the additive strategy in capturing the nuanced interplay of features pertinent to ABSA task.
@@ Appendix A Dataset Example
ARTS dataset employs three distinct strategies to rigorously test the model’s robustness: RevTgt is to generate sentences that reverse the original sentiment of the target aspect. RevNon is to change the target sentiment. AddDiff investigate if adding more nontarget aspects can confuse the model. We provide concrete instances of how each strategy is applied to manipulate aspect sentiment within the dataset in Table 6. Detailed statistics of the test sets are provided in Table 5.
@@ Appendix C ChatGPT Prompt
We conduct 3-shot prompting experiments on the ARTS datasets following Wang et al. (2023). We set the decoding temperature as 0 to increase ChatGPT’s determinism. The result is presented in Table 7.

### v2

@@ 3.4 Training and Inference
The loss component $\mathcal{L}_{K}$ corresponds to the cross-entropy loss calculated from the predictions of $\Psi(\zeta_{a},\zeta_{r^{{}^{\prime}}},\zeta_{k})$, as defined in Eq. (14). Similarly, the aspect-only and debiased review-only losses are denoted as $\mathcal{L}_{A}$ and $\mathcal{L}_{R}$ respectively.
@@ 4 Experiments
Table 1: We retrained BERT, RoBERTa as fair baselines ensuring that comparisons are made under similar training settings, which is crucial for validating DINER’s superior performance.
@@ 4.1 Datasets
We conduct training on the original SemEval 2014 Laptop and Restaurant datasets Pontiki et al. (2014), and perform testing on the ARTS datasets, as introduced by  Xing et al. (2020), to assess the efficacy of the proposed method. Detailed information about the ARTS datasets is shown in Appendix 6.
@@ 5 Result and Analysis
Table 1 presents a detailed comparison of various models’ performance for laptop and restaurant domains of the ARTS datasets, focusing on three key evaluation metrics: Acc., F1-score, and ARS.
@@
We evaluate DINER in two backbones: BERT-based and RoBERTa-based. These configurations are set to evaluate the effectiveness of DINER when integrated with different encoder-only PLMs. And DINER based on RoBERTa tends to outperform its BERT counterparts, which may be attributed to RoBERTa’s more robust pre-training on a larger and more diverse corpus, leading to better generalization capabilities Liu et al. (2019b). The results are compelling, showing that DINER(RoBERTa-based) model achieves the state-of-the-art performance across all metrics in both the laptop and restaurant domains, with a notable Acc. of 76.51 and 82.46, F1-scores of 73.27 and 76.92, and ARS of 59.40 and 64.02, respectively. DINER(RoBERTa-based) demonstrates superior performance in the Laptop domain, outpacing the baseline RoBERTa by margins of 1.55, 1.11, and 3.13 in terms of Acc., F1-score, and ARS metrics, respectively. In the Restaurant domain, the model further extends its lead, achieving improvements of 3.20, 6.45, and 4.06 in the same metrics. Similarly, the DINER(BERT-based) exhibits empirical enhancements.
@@ 5.1 More Detailed Result
RevTgt RevNon AddDiff Original
Laptop Vanilla 62.45 85.93 76.33 80.41
DINER 65.02($\uparrow$ 4.12%) 86.67($\uparrow$ 0.86%) 78.06($\uparrow$ 2.27%) 81.19($\uparrow$ 0.97%)
Restaurant Vanilla 64.06 82.66 83.48 85.18
DINER 70.69($\uparrow$ 10.35%) 83.56($\uparrow$ 1.08%) 86.07($\uparrow$ 3.10%) 87.32($\uparrow$ 2.51%)
Table 2: We use RoBERTa as the backbone. Vanilla refers to RoBERTa in Table 1. We compare the Acc. on the Vanilla and our DINER framework. We also calculate the change of accuracy.
We list in detail the performance of each model on the aforementioned three subsets of the ARTS datasets in Table 2. In the Laptop domain, the baseline model RoBERTa exhibits a RevTgt accuracy of 62.45, as RevTgt is the most challenging subset. It requires the model to pay precise attention to the target sentiment words. In contrast, the DINER framework significantly enhances this metric to 65.02, marking a 4.12% increment. Similarly, for RevNon and AddDiff, DINER outperforms the Vanilla baseline with modest improvements of 0.86% and 2.27%, respectively. The Restaurant domain further underscores the efficacy of the DINER framework, where a remarkable 10.35% improvement is observed in the RevTgt task, elevating the accuracy from 64.06 to 70.69. The framework also exhibits gains in RevNon and AddDiff tasks by 1.08% and 3.10%, respectively. The significant improvement observed in the restaurant domain underscores the effectiveness of our methods, particularly given the inherently challenging nature of the test set data in this domain, as highlighted by Xing et al. (2020). Specifically, the more challenging the dataset, the greater the improvement our framework offers. Interestingly, our method also gives a slight improvement on the Original test set, illustrating the fact that we have also de-biased the robust test data on the Original test set.
5.2 Effects of the Two Branches in DINER
Methods Laptop Restaurant
Vanilla 74.96 79.26
$R\rightarrow L$ branch
   + Causal Intervention(NWGM) 75.44 81.02
   + Causal Intervention(IPW) 75.50 81.19
    +$T\!D\!E$ 75.92 81.78
$A\rightarrow L$ branch
   + Counterfactual Inference 75.23 80.51
DINER 76.51 82.46
For the $R\rightarrow L$ branch, NWGM Xu et al. (2015) yields a marginal improvement in accuracy across both domains. The method of IPW Pearl (2009) further enhances performance, suggesting the efficacy of backdoor adjustment intervention in the methods, and IPW has a more precise approximation compared to NWGM Xu et al. (2015). We further debias context based on $T\!D\!E$, as described in Section 3.2, and performance is further enhanced upon the application of Counterfactual Reasoning.
5.3 Impact of Different Fusion Strategies
@@
Type Examples(Target Aspect: food) Gold Baseline DINER
Original The food is top notch, the service is attentive, and the atmosphere is great. Positive Positive Positive
RevTgt The food is nasty, but the service is attentive, and the atmosphere is great. Negative Negative Negative
RevNon The food is top notch, the service is heedless, but the atmosphere is not great. Positive Negative Positive
AddDiff The food is top notch, the service is attentive, and the atmosphere is great, but music is too heavy, waiters is angry and staff is arrogant. Positive Negative Positive
Table 5: Examples of case study. The corresponding gold labels and the predictions for each example are presented.
@@
$\left\{\begin{array}[]{l}\text{{MUL}-{Vanilla}}:L_{a,r^{{}^{\prime}},k}=\zeta_{a}\cdot\zeta_{r^{{}^{\prime}}}\cdot\zeta_{k},\\ \text{{MUL}-{sigmoid}}:L_{a,r^{{}^{\prime}},k}=\zeta_{k}\cdot\sigma(\zeta_{a})\cdot\sigma(\zeta_{r^{{}^{\prime}}}),\\ \text{{MUL}-$\tanh$}:L_{a,r^{{}^{\prime}},k}=\zeta_{k}\cdot\tanh(\zeta_{a})\cdot\tanh(\zeta_{r^{{}^{\prime}}}),\\ \text{{SUM}-{Vanilla}}:L_{a,r^{{}^{\prime}},k}=\zeta_{a}+\zeta_{r^{{}^{\prime}}}+\zeta_{k},\\ \text{{SUM}-{sigmoid}}:L_{a,r^{{}^{\prime}},k}=\zeta_{k}+\sigma(\zeta_{a})+\sigma(\zeta_{r^{{}^{\prime}}}),\\ \text{{SUM}-$\tanh$}:L_{a,r^{{}^{\prime}},k}=\zeta_{k}+\tanh(\zeta_{a})+\tanh(\zeta_{r^{{}^{\prime}}})\end{array}\right.$ (16)
The Acc. performance of six distinct different fusion strategies are reported in Table 4. From the table, we can find that the MUL fusion, regardless of the activation function, consistently underperforms in comparison to its SUM counterparts. Apparently, SUM fusion strategies are more stable and robust, and more suitable for the ASBA task. The superior performance of SUM fusion strategies, particularly with the $\tanh$ activation, underscores the effectiveness of the additive strategy in capturing the nuanced interplay of features pertinent to the ABSA task.
5.4 Case Study
To demonstrate the efficacy of the proposed method, we present a case study featuring a sample and its three adversarial variants in Table 5. We compare our proposed method based on RoBERTa with the baseline RoBERTa.
From the table, the results clearly demonstrate that our method DINER, exhibits enhanced robustness compared to the baseline approach. Specifically, Original and RevTgt types, where either no changes or direct negative changes were made to the targeted aspect, both methods perform equally well.
However, the distinction in performance is evident in more complex adversarial examples. In the RevNon type, where distractors are introduced in non-target aspects (e.g., service and atmosphere), the baseline fails to maintain its accuracy, misclassifying the overall sentiment as Negative. In contrast, DINER successfully recognizes the sentiment as Positive, reflecting its ability to isolate the influence of perturbations to non-target aspects. The AddDiff type further complicates the scenario by adding multiple negative aspects unrelated to the target. Despite these challenges, DINER continues to accurately assess the sentiment towards the food as Positive, whereas the baseline erroneously shifts to a Negative prediction.
The resilience of our method to adversarial conditions suggests it is well-suited for real-world environments where reliable sentiment analysis is crucial.
@@ 6 Conclusion
Acknowledgement
The authors would like to thank the anonymous reviewers for their insightful comments. This work is funded by the National Natural Science Foundation of China (62176053). This work is supported by the Big Data Computing Center of Southeast University.
@@ Appendix A Dataset Example
ARTS datasets employ three distinct strategies to rigorously test the model’s robustness: RevTgt is to generate sentences that reverse the original sentiment of the target aspect. RevNon is to change the target sentiment. AddDiff investigate if adding more nontarget aspects can confuse the model. We provide concrete instances of how each strategy is applied to manipulate aspect sentiment within the dataset in Table 7. Detailed statistics of the test sets are provided in Table 6.
@@ Appendix C ChatGPT Prompt
We conduct 3-shot prompting experiments on the ARTS datasets following Wang et al. (2023). We set the decoding temperature as 0 to increase ChatGPT’s determinism. The prompts are presented in Table 8.